# TraceLens: Question-Driven Debugging for Taint Flow Understanding

### v1

@@ III-A CodeQL
Figure 8: CodeQL’s tree-based view shows each warning one by one and does not support inquiry-based sensemaking of multiple taint flows and the impact of user choices and 3rd-party library models.
III-B Soufflé and Logic Query Implementation
TraceLens converts the CodeQL taint-analysis results into a set of facts that can be analyzed by Soufflé’s Datalog engine. In this factbase, each node, edge, and taint-related annotation (e.g., source, sink, and library API) is expressed as a logical predicate. Once the output of a single CodeQL query is transformed into these facts, TraceLens can quickly re-evaluate multiple “secondary” queries, such as those in Table I, without needing to re-run CodeQL’s underlying taint analysis itself.
This arrangement is particularly useful for tasks like missing-flow analysis (WhyNotFlow) or identifying global impact. Instead of incurring repeated analysis times of 10 to 14 seconds (excluding the compile time of the query) in CodeQL for each reconfiguration of the library model, we rely on Soufflé’s Datalog engine [15] to query the precomputed facts in the order of 5 seconds—a speed-up that supports interactive debugging.
@@ III-B Soufflé and Logic Query Implementation
Once loaded into Soufflé, the questions in Table I are encoded as logic queries shown in Figures 2 and 3. For example, WhyFlow query uses transitive closure over edge to find whether data can reach a sink from a source, intersecting with library_flow to highlight which third party library’s APIs (orange nodes) appear along the path. WhyNotFlow similarly checks for “broken” transitive paths and identifies library_flow edges that act as sanitizers, effectively terminating any flow.
@@
III-C TraceLens’s User Interface
To support sensemaking, key design elements in TraceLens include:
Color-Coding and Visual Clarity.
Nodes in the flow graph are color-coded: green for sources, red for sinks, orange for external APIs or libraries, and blue for other intermediate nodes.
@@ III-C TraceLens’s User Interface
TraceLens overlays the taint flows onto a single graph. Solid edges represent dataflow steps. Dashed edges indicate “plausible flows” currently unreported by the analyzer, but can be reported under a different configuration. Each flow can be expanded to narrow down onto suspicious segments.
@@
Clicking on a node opens the corresponding code snippet in the user’s IDE for easy reference of the source code. TraceLens includes hover popups showing details such as fully-qualified names of identifiers referenced by the intermediate steps along the flow.
@@
TraceLens supports multiple layout algorithms (e.g., breadth-first, concentric). Users can reorganize the graph for improved clarity. A GlobalImpact query resizes API nodes by their number of occurrences on different flows.
IV User Study
To evaluate TraceLens’s usefulness in sensemaking taint flows, we designed a within-subject study. We assess how users can reason about the impact of user configurations on multiple taint flows, including third-party taint analysis models. We use CodeQL’s Visual Studio Code plugin (CodeQL Visualizer in short), as the baseline.
@@ IV User Study
Participants were asked to answer eight sensemaking questions about taint analysis results and how the configuration of third-party libraries impacts taint flows. Each user study task consists of eight questions in Table III. The first six questions centered on why, why-not, and what-if questions about taint flows and the remaining two centered on quantifying taint flows.
@@
  1. 1.
    How much does TraceLens improve the participants’ ability to answer questions about the configuration’s impact on taint flows?
  2. 2.
    How does TraceLens influence cognitive load and user confidence in sensemaking taint flows?
  3. 3.
    What are the participants’ perceptions of TraceLens’s usability and functionality in enhancing their workflow?
IV-A Study Protocol
@@
Table II: Statistics of Taint Analysis Facts for Apache Dubbo.
The study task is based on 383 taint warnings generated on Apache Dubbo with the CodeQL query shown in Listing 1 [16]. The resulting facts from this query are presented in Table II.
Listing 1: CodeQL Taint Analysis Query [16].
@@ IV-A2 Participants
While between-subject user studies require a large number of participants to account for variations among individual participants, within-subject user studies minimize variability, as each participant uses both tools following a different order. The order of tool usage and task assignment is randomized and counterbalanced [17, 18]. Within-subject user studies with 8 to 16 participants are standard practice in both software engineering [19, 20, 21] and HCI research [22, 23, 24].
@@
While most of our study’s participants are students, prior work found that the findings based on student participants often generalize to a broader population as students have comparable performance to professionals in security-related tasks [25, 26, 27, 28]. Other studies [29, 30, 31] also emphasize that students can provide meaningful feedback when their expertise aligns with the tool’s target audience. Since TraceLens is designed for end-user debugging, the use of student participants is methodologically sound. We also conducted a case study with two security professional participants, discussed in Section LABEL:casestudy. Also, the goal of TraceLens is to help end-users who are not necessarily security experts. TraceLens helps them to sensemake taint analysis results and debug them with our interrogative queries.
IV-A3 Protocol
Each participant took part in a 1.5-hour session. The study involved using both TraceLens and the baseline CodeQL visualizer and the two tasks. The order of assigned tool (TraceLens first vs. CodeQL first) and the assigned task (Problem Set #1 and Problem Set #2) was counterbalanced across participants through random assignment. We gave a 3-minute pre-study survey to collect background information, followed by a tutorial.
Next, each participant proceeded to the two tasks, each lasting 20 minutes. Each task had 8 questions that participants were tasked to answer. They had the option of skipping over questions and ending the task early. After each task, participants filled out post-task survey, which included the NASA TLX (5 minutes total). In the final 12-minute post-study survey, we asked participants to compare TraceLens and CodeQL, and rate TraceLens’s features.
IV-A4 Tutorial
We conducted a tutorial session to introduce participants to the functionality of both TraceLens and the CodeQL VSCode plugin. The tutorials were structured around realistic dataflow scenarios. For TraceLens, we guided users through all six query types. This involved selecting relevant sources, sinks, or API calls from dropdown menus and interpreting the graphical output rendered by TraceLens. Participants were encouraged to hover over nodes to view fully qualified names, click on nodes to navigate directly to code locations, and adjust the layout using built-in options. Through these interactions, participants familiarized themselves with TraceLens’s features.
@@ IV-A4 Tutorial
IV-A5 Tasks
Each task had a different set of questions (Problem Set #1 and Problem Set #2). The questions and their answers are shown in Table III. To minimize learning effects, we balanced the order of tools that participants used first. Each task contained eight questions about ’why’, ’why-not’, ’what-if’, shown in Table I and the other two about quantitative aspects of taint flows (e.g., how many taint paths exist or how many third-party APIs appear in the taint paths). The questions were multiple-choice questions, some had multiple answers. We granted partial credit if participants selected some correct answers, but missed or added incorrect choices.
@@ IV-A5 Tasks
Table III: Study Tasks: Program Sets #1 and #2 with their correct answers. Participants were assigned one of the two tasks to complete with CodeQL visualizer and the other with TraceLens.
IV-A6 Criteria for Study Tasks
@@
After completing each task, participants completed a brief post-task questionnaire, which included the NASA TLX [32] (e.g., mental demand, time pressure, perceived success, effort, and frustration) and described their strategies to find answers to the questions. At the end of the session, participants rated user-interface features, confidence, and ease of use.
V Results
VI Discussion
Listing 2: A divergent path query flagging branch points where flow diverges to multiple targets. Using the raw three-arity edge(id,src,dst) input, we declare branch(n) for each $n$ that appears as src in more than one edge(_,n,_) fact.
@@ VI Discussion
Participants using TraceLens achieved better accuracy, especially when analyzing multiple flows. In particular, the AffectedSinks and GlobalImpact queries require the analysis of multiple or interconnected flows. Surprisingly, participants were able to answer questions related to “why-not” using both tools. In contrast, participants using the CodeQL visualizer responses frequently had empty submissions due to lack of time, suggesting viewing results one by one a list view is time-consuming. This shows that an interactive, inquiry-based debugger of taint analysis warnings has the potential to improve sensemaking for large result sets, which is known to hinder adoption [10].
@@
While participants found TraceLens “less cumbersome” (P2) for analysis involving multiple flows, they offered suggestions for improving TraceLens. They recommended adding an on-screen legend to clarify colors and dotted edges, and using distinct shapes for different node types (e.g., sources versus sinks) to avoid visual confusion. Additional interactive features, such as “click-to-edit” functionality for enabling or disabling sanitizers were also proposed. This suggests that participants would appreciate real-time feedback through interaction on the graph.
@@
While TraceLens supports six template questions from Table I, adding more template queries is easy since TraceLens’s template questions are Datalog-style queries. As a case study, adding a new template query called ’divergent path’ query to TraceLens (Listing 2) took under ten minutes. This query identifies the branch points, which are locations where the flow branches into multiple targets. Note that TraceLens’s queries supports any taint analysis results from CodeQL. More extension queries are available in our replication package (https://github.com/tracelens/TraceLens/tree/main/data/extension_queries).
@@
TraceLens supports interrogative and speculative debugging built on a rich history of using logic programming for software comprehension [41, 42, 43, 44, 45, 46]. Soufflé is the state of the art logic query engine. TraceLens does not compete with underlying taint analysis engines but instead emphasizes sensemaking support. We pay a one-time cost of fact extraction from CodeQL’s result to support TraceLens’s interactive features.
@@
A threat to validity is the size of our user study. While our user study involved only 12 participants, it employs a within-subject crossover design, thereby reducing variance and requiring fewer participants [17, 47]. While our study evaluated TraceLens only against a single baseline (CodeQL) and and a single subject program (Apache Dubbo), our approach is not tied to CodeQL, and can generalize to any taint analysis.
VII Related Work
@@
Speculative What-If Analysis. Speculative execution allows the investigation of future or alternative actions developers may perform. Brun et al. explored its application for providing fix suggestions [48], as well as identifying conflicts during version merging [49]. Our work is the first to apply it to taint analysis for reasoning about permissible dataflows of sensitive information under different configuration of different dataflow connectivity.
@@
Templated Questions for asking questions. Developers often ask questions during variety of developer tasks [50]. In particular, developers need support for understanding code reachability [8], exploring how different vulnerabilities relate and finding similar ones [11], tracking intermediate states of analysis [51], reasoning about system-wide implications when making changes [11]. TraceLens is the first work providing templated questions for investigating and debugging the results of taint analysis.
@@
Modelling 3rd party libraries. Existing work [52, 53, 54] automatically infers models of third-party libraries. However, they do not guide users to understand the impact of these models. They do not support end-user debugging and do not allow end-users to ask questions about the impact of modelling choices. While Paralib [55] allows comparison between multiple choices of libraries, it does not allow reasoning about how incorrect models cause a mismatch between user expectations and the actual tool results.
@@
End-user debugging. End user debugging [56, 57, 58] is about the problem of investigating the root cause of tool outcomes, and has been applied to spreadsheet debugging, search engine configuration, etc. Similarly, TraceLens is an end-user debugger for reasoning how taint flows are affected by the configuration of sources, sinks, and models of third-party libraries.
VIII Conclusion
@@
We presented TraceLens, a tool for end-user debugging for taint analysis. Through speculative analysis, TraceLens allows developers to ask “why”, “why-not”, and “what-if” questions about the analysis configuration, and is able to visualize multiple, interconnected flows on a graphical view. TraceLens enables sensemaking of taint analysis and helps users identify root causes of unexpected flows or missing flows. Our user study confirms that TraceLens helps users provide 21% more correct answers to questions about the taint analysis, while experiencing a lower cognitive load. These results suggest that TraceLens’s inquiry-based approach empowers developers to analyze the outputs of taint analysis.
IX Data Availability
@@
The replication package and study’s data are available. at https://github.com/tracelens/TraceLens/tree/main/data/extension_queries.

### v2

@@ 3.1. CodeQL
Figure 8. CodeQL’s tree-based view shows each warning one by one and does not support inquiry-based sensemaking of multiple taint flows and the impact of user choices and 3rd-party library models.
3.2. Soufflé and Logic Query Implementation
WhyFlow converts the CodeQL taint-analysis results into a set of facts that can be analyzed by Soufflé’s Datalog engine. In this factbase, each node, edge, and taint-related annotation (e.g., source, sink, and library API) is expressed as a logical predicate. Once the output of a single CodeQL query is transformed into these facts, WhyFlow can quickly re-evaluate multiple “secondary” queries, such as those in Table 1, without needing to re-run CodeQL’s underlying taint analysis itself.
This arrangement is particularly useful for tasks like missing-flow analysis (WhyNotFlow) or identifying global impact. Instead of incurring repeated analysis times of 10 to 14 seconds (excluding the compile time of the query) in CodeQL for each reconfiguration of the library model, we rely on Soufflé’s Datalog engine (jordan2016souffle, ) to query the precomputed facts in the order of 5 seconds—a speed-up that supports interactive debugging.
@@ 3.2. Soufflé and Logic Query Implementation
Once loaded into Soufflé, the questions in Table 1 are encoded as logic queries shown in Figures 2 and 3. For example, WhyFlow query uses transitive closure over edge to find whether data can reach a sink from a source, intersecting with library_flow to highlight which third party library’s APIs (orange nodes) appear along the path. WhyNotFlow similarly checks for “broken” transitive paths and identifies library_flow edges that act as sanitizers, effectively terminating any flow.
@@
3.3. WhyFlow’s User Interface
To support sensemaking, key design elements in WhyFlow include:
Color‐Coding and Visual Clarity.
Nodes in the flow graph are color‐coded: green for sources, red for sinks, orange for external APIs or libraries, and blue for other intermediate nodes.
@@ 3.3. WhyFlow’s User Interface
WhyFlow overlays the taint flows onto a single graph. Solid edges represent dataflow steps. Dashed edges indicate “plausible flows” currently unreported by the analyzer, but can be reported under a different configuration. Each flow can be expanded to narrow down onto suspicious segments.
@@
Clicking on a node opens the corresponding code snippet in the user’s IDE for easy reference of the source code. WhyFlow includes hover popups showing details such as fully-qualified names of identifiers referenced by the intermediate steps along the flow.
@@
WhyFlow supports multiple layout algorithms (e.g., breadth‐first, concentric). Users can reorganize the graph for improved clarity. A GlobalImpact query resizes API nodes by their number of occurrences on different flows.
4. User Study
To evaluate WhyFlow’s usefulness in sensemaking taint flows, we designed a within-subject study. We assess how users can reason about the impact of user configurations on multiple taint flows, including third-party taint analysis models. We use CodeQL’s Visual Studio Code plugin (CodeQL Visualizer in short), as the baseline.
@@ 4. User Study
Participants were asked to answer eight sensemaking questions about taint analysis results and how the configuration of third-party libraries impacts taint flows. Each user study task consists of eight questions in Table 3. The first six questions centered on why, why-not, and what-if questions about taint flows and the remaining two centered on quantifying taint flows.
@@
  1. (1)
    How much does WhyFlow improve the participants’ ability to answer questions about the configuration’s impact on taint flows?
  2. (2)
    How does WhyFlow influence cognitive load and user confidence in sensemaking taint flows?
  3. (3)
    What are the participants’ perceptions of WhyFlow’s usability and functionality in enhancing their workflow?
@@
Table 2. Statistics of Taint Analysis Facts for Apache Dubbo.
4.1. Study Protocol
The study task is based on 383 taint warnings generated on Apache Dubbo with the CodeQL query shown in Listing 1 (codeql_externalapi_query, ). The resulting facts from this query are presented in Table 2.
Listing 1: CodeQL Taint Analysis Query (codeql_externalapi_query, ).
@@ 4.1.2. Participants
While between-subject user studies require a large number of participants to account for variations among individual participants, within-subject user studies minimize variability, as each participant uses both tools following a different order. The order of tool usage and task assignment is randomized and counterbalanced (vegas2015crossover, ; chasins_2021, ). Within-subject user studies with 8 to 16 participants are standard practice in both software engineering (garcia_2024, ; kang_2024, ; ganji_2023, ) and HCI research (horvath_2022, ; suh_2023, ; huh_2024, ).
@@
While most of our study’s participants are students, prior work found that the findings based on student participants often generalize to a broader population as students have comparable performance to professionals in security-related tasks (sandoval_2023, ; acar_2016, ; acar_2017, ; salman_2015, ). Other studies (naiakshina_2017, ; naiakshina_2020, ; ko_2015, ) also emphasize that students can provide meaningful feedback when their expertise aligns with the tool’s target audience. Since WhyFlow is designed for end-user debugging, the use of student participants is methodologically sound. We also conducted a case study with two security professional participants, discussed in Section LABEL:casestudy. Also, the goal of WhyFlow is to help end-users who are not necessarily security experts. WhyFlow helps them to sensemake taint analysis results and debug them with our interrogative queries.
4.1.3. Protocol
Each participant took part in a 1.5-hour session. The study involved using both WhyFlow and the baseline CodeQL visualizer and the two tasks. The order of assigned tool (WhyFlow first vs. CodeQL first) and the assigned task (Problem Set #1 and Problem Set #2) was counterbalanced across participants through random assignment. We gave a 3-minute pre-study survey to collect background information, followed by a tutorial.
Next, each participant proceeded to the two tasks, each lasting 20 minutes. Each task had 8 questions that participants were tasked to answer. They had the option of skipping over questions and ending the task early. After each task, participants filled out post-task survey, which included the NASA TLX (5 minutes total). In the final 12-minute post-study survey, we asked participants to compare WhyFlow and CodeQL, and rate WhyFlow’s features.
4.1.4. Tutorial
We conducted a tutorial session to introduce participants to the functionality of both WhyFlow and the CodeQL VSCode plugin. The tutorials were structured around realistic dataflow scenarios. For WhyFlow, we guided users through all six query types. This involved selecting relevant sources, sinks, or API calls from dropdown menus and interpreting the graphical output rendered by WhyFlow. Participants were encouraged to hover over nodes to view fully qualified names, click on nodes to navigate directly to code locations, and adjust the layout using built-in options. Through these interactions, participants familiarized themselves with WhyFlow’s features.
@@ 4.1.4. Tutorial
4.1.5. Tasks
Each task had a different set of questions (Problem Set #1 and Problem Set #2). The questions and their answers are shown in Table 3. To minimize learning effects, we balanced the order of tools that participants used first. Each task contained eight questions about ’why’, ’why-not’, ’what-if’, shown in Table 1 and the other two about quantitative aspects of taint flows (e.g., how many taint paths exist or how many third-party APIs appear in the taint paths). The questions were multiple-choice questions, some had multiple answers. We granted partial credit if participants selected some correct answers, but missed or added incorrect choices.
@@ 4.1.5. Tasks
Table 3. Study Tasks: Program Sets #1 and #2 with their correct answers. Participants were assigned one of the two tasks to complete with CodeQL visualizer and the other with WhyFlow.
4.1.6. Criteria for Study Tasks
@@
After completing each task, participants completed a brief post-task questionnaire, which included the NASA TLX (hart1988development, ) (e.g., mental demand, time pressure, perceived success, effort, and frustration) and described their strategies to find answers to the questions. At the end of the session, participants rated user-interface features, confidence, and ease of use.
5. Results
6. Discussion
Listing 2: A divergent path query flagging branch points where flow diverges to multiple targets. Using the raw three‐arity edge(id,src,dst) input, we declare branch(n) for each $n$ that appears as src in more than one edge(_,n,_) fact.
@@ 6. Discussion
Participants using WhyFlow achieved better accuracy, especially when analyzing multiple flows. In particular, the AffectedSinks and GlobalImpact queries require the analysis of multiple or interconnected flows. Surprisingly, participants were able to answer questions related to “why-not” using both tools. In contrast, participants using the CodeQL visualizer responses frequently had empty submissions due to lack of time, suggesting viewing results one by one a list view is time-consuming. This shows that an interactive, inquiry-based debugger of taint analysis warnings has the potential to improve sensemaking for large result sets, which is known to hinder adoption (johnson2013don, ).
@@
While participants found WhyFlow “less cumbersome” (P2) for analysis involving multiple flows, they offered suggestions for improving WhyFlow. They recommended adding an on-screen legend to clarify colors and dotted edges, and using distinct shapes for different node types (e.g., sources versus sinks) to avoid visual confusion. Additional interactive features, such as “click-to-edit” functionality for enabling or disabling sanitizers were also proposed. This suggests that participants would appreciate real-time feedback through interaction on the graph.
@@
While WhyFlow supports six template questions from Table 1, adding more template queries is easy since WhyFlow’s template questions are Datalog-style queries. As a case study, adding a new template query called ’divergent path’ query to WhyFlow (Listing 2) took under ten minutes. This query identifies the branch points, which are locations where the flow branches into multiple targets. Note that WhyFlow’s queries supports any taint analysis results from CodeQL. More extension queries are available in our replication package (https://github.com/whyflowtaint/WhyFlow/tree/main/data/extension_queries).
@@
WhyFlow supports interrogative and speculative debugging built on a rich history of using logic programming for software comprehension (holt_1998, ; mens_2002, ; hajiyev_2005, ; eichberg_2008, ; guéhéneuc_2008, ; tourwe_2003, ). Soufflé is the state of the art logic query engine. WhyFlow does not compete with underlying taint analysis engines but instead emphasizes sensemaking support. We pay a one-time cost of fact extraction from CodeQL’s result to support WhyFlow’s interactive features.
@@
A threat to validity is the size of our user study. While our user study involved only 12 participants, it employs a within-subject crossover design, thereby reducing variance and requiring fewer participants (vegas2015crossover, ; Gergle_Tan_2014, ). While our study evaluated WhyFlow only against a single baseline (CodeQL) and and a single subject program (Apache Dubbo), our approach is not tied to CodeQL, and can generalize to any taint analysis.
7. Related Work
@@
Speculative What-If Analysis. Speculative execution allows the investigation of future or alternative actions developers may perform. Brun et al. explored its application for providing fix suggestions (brun2010speculative, ), as well as identifying conflicts during version merging (brun2011proactive, ). Our work is the first to apply it to taint analysis for reasoning about permissible dataflows of sensitive information under different configuration of different dataflow connectivity.
@@
Templated Questions for asking questions. Developers often ask questions during variety of developer tasks (begel2014analyze, ). In particular, developers need support for understanding code reachability (ko2004designing, ), exploring how different vulnerabilities relate and finding similar ones (smith2015questions, ), tracking intermediate states of analysis (do2018debugging, ), reasoning about system-wide implications when making changes (smith2015questions, ). WhyFlow is the first work providing templated questions for investigating and debugging the results of taint analysis.
@@
Modelling 3rd party libraries. Existing work (li2024llm, ; piskachev2019swan_assist, ; Chiang_Li_Zhou_Banerjee_Schaef_Lyu_Nguyen_Tripp_2024, ) automatically infers models of third-party libraries. However, they do not guide users to understand the impact of these models. They do not support end-user debugging and do not allow end-users to ask questions about the impact of modelling choices. While Paralib (yan2022concept, ) allows comparison between multiple choices of libraries, it does not allow reasoning about how incorrect models cause a mismatch between user expectations and the actual tool results.
@@
End-user debugging. End user debugging (burnett2004end, ; kissinger2006supporting, ; grigoreanu2012end, ) is about the problem of investigating the root cause of tool outcomes, and has been applied to spreadsheet debugging, search engine configuration, etc. Similarly, WhyFlow is an end-user debugger for reasoning how taint flows are affected by the configuration of sources, sinks, and models of third-party libraries.
8. Conclusion
@@
We presented WhyFlow, a tool for end-user debugging for taint analysis. Through speculative analysis, WhyFlow allows developers to ask “why”, “why-not”, and “what-if” questions about the analysis configuration, and is able to visualize multiple, interconnected flows on a graphical view. WhyFlow enables sensemaking of taint analysis and helps users identify root causes of unexpected flows or missing flows. Our user study confirms that WhyFlow helps users provide 21% more correct answers to questions about the taint analysis, while experiencing a lower cognitive load. These results suggest that WhyFlow’s inquiry-based approach empowers developers to analyze the outputs of taint analysis.
9. Data Availability
@@
The replication package and study’s data are available. at https://github.com/whyflowtaint/WhyFlow.